|
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
SPB
SQS
XO
ToBeMigrated2XWiki
ToBeMigrated2GitlabPages
ToBeMigratedByGroup
Migrated2Zulip
Archived2PDF
not2MigrateOrArchive
|
|
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB Migrated2Zulip
|
|
SASE1 Status and SASE3 Readiness
Technical integration and coordination
|
| Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
|
Message ID: 30
Entry time: 08 Apr 2019, 10:48
|
| Author: |
Adriano Violante |
| Location: |
General |
| Severity: |
Information |
| Subject: |
Report after dispatch meeting 09.30 |
|
|
SPB/SFX
- 5Hz mode tests started (SPB/SQS) and partially successful. A detailed report will be sent around. 2 cameras in SPB and digitizers in SQS were tested. Cameras OK. Digitizers data recordin has some issues being fixed (train id related). Next week on Thursday a test with beam is foreseen. ==> Update from AS:The whole chain, Digitizer, Karabo, DAQ worked as expected.
- CSS being commissioned
- AGIPD chiller: Julabo is back. Loop is being updated. Detector interlocks also being updated.
FXE
SQS
- Loop 8 and 11 deployed
- Triggers to be re-checked (BF+TB)
- Configuration of motors is being finalized.
- Network drive in the control network. Solution will be discussed with ITDM (group access to the DESY cloud?)
SCS
- XS5000 controllers in the laser room were not safe and need to be reworked. Technically the issue has been solved. SCS reported some unhappiness because of unclear procedures (responsibility, Elektrofachkraefte...). AV will discuss with SRP and UBr, TS and ask SRP to make a statement.
- Shutter in the experimant hutch will stay open the entire day to allow XRO to continue the commisisoning of the mirror and mono in the tunnels.
MID
- Epix - essentially ready for experiments this week.
- 2nd epix module being tested.
- Epix channel power issue (digital/analog voltage issue). Being investigated. It is not a showstopper
HED
- Loop 5 will be commisisoned today. Loop 7 will follow (much bigger, in-vacuum components). The goal is to finish the commissioning of these 2 loops by the end of the week.
OPT Lasers
XRO
- SA3 commissioning during the week end (Mirror+mono). Will continue today.
- Procedure to re start the loops in the tunnels has been discussed and modified to increase the transparency. PG will write an email summarizing the new procedure to the GLs
XPD
VAC
DET
- HED epix and jungfrau calibration routines will be installed tomorrow.
- FXE: LPD charge injection scan has to be re-scheduled.
AE
ITDM
EET
CAS
SRP
TS
AoB |